OpenStudy (anonymous):

Hypotenuse = 10 Long =

OpenStudy (jadeishere):

well 10^2 = a^2 + b^2 are you trying to solve for the legs?

OpenStudy (anonymous):

yes. I have to put it into an equation

OpenStudy (anonymous):

Mason owns a collection of hardback books in a series about American history. Each book has dimensions as shown below. There are 7 books in the series, stacked side by side. What is the volume of the entire collection? a) one hundred sixty-one cubic inches b) eight hundred forty cubic inches c) nine hundred twenty-four cubic inches d) nine hundred sixty cubic inches

OpenStudy (anonymous):

help

OpenStudy (jadeishere):

well 10^2 = 100 100 = a^2 + b^2 so we know both a and b are below 10

OpenStudy (anonymous):

AD = 6 DB = 2 AC =

OpenStudy (anonymous):

watt about this question

OpenStudy (jadeishere):

Well, now sure how to solve that @Night-Watcher

OpenStudy (anonymous):

plzz

OpenStudy (anonymous):

A rectangular prism, representing a book, has a LENGTH of 1 inch, a width of ten inches, and a height of twelve inches. 1×10 is the base So, base times height, 10×12 I think the final answer might be 120.